### v1

@@ 5 Evaluation
  • Venn speeds up the average JCT by up to 1.88X without affecting the model accuracy, compared to the state-of-the-art across various real-world FL workloads (5.2).
  • Venn outperforms its counterparts through intelligent job scheduling and device-to-job matching using different design components (5.3).
  • Venn ’s benefits are robust under a wide range of FL workloads and environment setups (5.5).
@@ 5.1 Experiment Setup
To rigorously evaluate Venn, we employ a two-pronged approach. First, we have developed a high-fidelity simulator that replays client and job traces, effectively emulating the dynamics of the scheduling environment for large-scale evaluations. Second, we deploy real FL systems to execute actual FL jobs at a smaller scale of devices.
FL Resources:
@@
To faithfully emulate heterogeneous device runtimes, network throughput, and availability, we use device traces from FedScale Lai et al. (2021a) and AI Benchmark Ignatov et al. (2019), as depicted in Figure 2 and Figure 7(a). Each unique device trace is limited to one FL job per day for realism.
FL Jobs:
@@
Our focus is on synchronous FL jobs Yang et al. (2018), where each successful training round requires a minimum of 80% target participants to report back within a deadline, which is set to be 5min - 15min depending on the round demand. To assess the generalizability of Venn, we curate a diverse set of FL jobs drawn from diverse applications Xu et al. (2023); Yang et al. (2018); Hard et al. (2019); Ramaswamy et al. (2019); Warden (2018); Sadilek et al. (2021), whose resource demand is depicted in Figure 7(b). In the real FL experiment, each job aims to train a ResNet-18 He et al. (2016) and MobileNet-V2 Sandler et al. (2018) on FEMNIST dataset.
Workloads:
@@
Our evaluation includes five workload scenarios that sample differently from the job trace in Figure 7(b) to rigorously evaluate Venn’s performance. Even: Sampled from all jobs, which is the default trace. Small: Uniformly sampled only from jobs with below-average total demand. Large: Uniformly sampled only from jobs with above-average total demand. Low: Uniformly sampled only from jobs with below-average demand per round. High: Uniformly sampled only from jobs with above-average demand per round.
@@
Devices are stratified into four specifications based on their CPU and memory capacities (Figure 7(a)) to create various resource contention patterns where the eligible resource set for one job may overlap, contain, or be within the eligible resource set of other jobs. By default, each job is randomly mapped to one device specification among General resources, Compute-Rich resources, Memory-Rich resources, High-Performance resources. Jobs arrive via a Poisson process with a 30-min average inter-arrival. Default simulation and real-world workloads contain 50 and 20 jobs, respectively.
Baselines
@@
We compare Venn with FIFO, SRSF, and an optimized random matching. Both Venn and SRSF are agnostic to the total FL job rounds. Random matching algorithm is supposed to match one eligible job to each device at random. However, to reduce its round abortion rate under contention, we optimize it to schedule job requests in a randomized order, thereby setting a more challenging baseline. Note that we only run Venn with the starvation prevention strategy in Section 5.5.
Metrics:
@@
Our primary performance metrics include the average job completion time (JCT). Note that while Venn does not explicitly optimize for FL job accuracy, it does not adversely affect it either.
@@
Table 1: Summary of improvements on average JCT over random matching on different FL workloads.
@@ 5.2 End-to-End Performance
Venn achieves better average JCT Improvement. We assess the performance of different scheduling algorithms by evaluating its performance over different workloads. We report the average JCT speed-up for each scheduling algorithms compared to the random scheduling in Table 1. We observe that our scheduling algorithm consistently provides stable improvements in the average JCT across various workloads, which underscores the robustness of Venn.
Venn achieves faster convergence without affecting accuracy. We report the final model test accuracy of FL jobs under different schedules with the help of our FL system at a smaller experiment scale. As shown in Figure 10, we observe that Venn does not affect the final model test accuracy but speeds up the overall convergence process.
Venn has negligible overhead. We emulated a large number of FL jobs and groups to evaluate the scheduler’s scalability. Our results in Figure 10 demonstrate that the latency incurred by one-time triggering for scheduling and matching remains low, even with a substantial increase in job and group numbers. This can be attributed to its time complexity $\max(O(m\log m),O(n^{2}))$, where $m$, $n$ are the numbers of ongoing jobs and job groups, respectively.
Figure 9: Venn does not affect the average test accuracy.
@@ 5.3 Performance Breakdown
Table 3: Breakdown of average JCT Improvement across jobs that ask for different resource requirements: General resources, Compute-rich resources, Memory-rich resources and High-performance resources. Venn benefits more on jobs that ask for scarcer resources.
To analyze the impact of Venn on different types of jobs, we break down jobs by their total demands and device requirements (Figure 7(a)), and then analyze the average JCT improvement for each type. Table 2 and Table 3 quantifies how Venn improves average JCT across varying total demands (25th, 50th, 75th percentiles) and eligibility types (General, Compute-rich, Memory rich, High-performance ). Notably, jobs that have smaller total demands and ask for scarce resources benefit the most from Venn.
5.4 Case Study on Biased Workload
@@ 5.5 Ablation Study
Impact of number of jobs:
We evaluate Venn under various scenarios with different numbers of FL jobs arriving over time. As the number of jobs increases, resource contention becomes more pronounced, highlighting the importance of efficient scheduling under such conditions. We present the average JCT speed-up with different numbers of jobs in even workload to demonstrate the effectiveness of our algorithm. As shown in Figure 13, Venn consistently provides improvement across various numbers of jobs, with its benefits becoming more pronounced as the number of jobs increases.
Impact of number of tiers:
@@
Impact of fairness knob:
We incorporated a fairness knob ($\epsilon$) to strike a balance between performance and fairness. We conducted an evaluation using various values of $\epsilon$, plotting the average JCT speed-up against these values in Figure 13(a). The results demonstrate that, as $\epsilon$ increases, the JCT speed-up tends to decrease. As shown in Figure 13(b), the percentage of jobs that meet the fair-share JCT increases with the $\epsilon$, where ϵ=italic-$\epsilon=$2 gives 69% of jobs receive their fair-share JCT. This observation highlights the trade-off between performance and fairness within our FL resource scheduling algorithm, which can be fine-tuned by adjusting the value of $\epsilon$.
6 Related Works
Cluster Resource Manager. There are many cluster resource managers that schedule heterogeneous resources with constraints Thinakaran et al. (2017); Ghodsi et al. (2013); Narayanan et al. (2020). However, they are mostly designed for data center and fail to capture the level of availability, heterogeneity and scale of FL resources, nor do they consider both scheduling delay and response collection time in the JCT. Existing ML cluster resource managers mainly focus on managing the stable data-center resources like GPU and CPU (Xiao et al., 2018; Mohan et al., 2022) in order to improve JCT, utilization and fairness (Peng et al., 2018; Chaudhary et al., 2020). Some research delves into GPU-specific optimizations Gu et al. (2019); Yu et al. (2021); Hwang et al. (2021), while others co-design resource managers with the specific characteristics of ML workloads Qiao et al. (2021); You et al. (2023). However, these ML cluster resource managers are not directly applicable to the more uncertain and heterogeneous FL resources, nor to FL job characteristics. Moreover, we want to note that Venn is complementary to these in-cluster schedulers. For example, when the number of rounds of FL jobs is not available, we can leverage Least-Attained Service Gu et al. (2019) to performed Venn scheduling within each Tiresias-determined job queue.
FL Client Selector. Several recent works have studied client selection at the single FL job level. In addition to enforce device requirements including software version, hardware capacity and data quality, they further cherry-pick clients based on their state, system and statistical utility (Lai et al., 2021b; Abdelmoniem et al., 2023; Balakrishnan et al., 2022; Jee Cho et al., 2022; Chai et al., 2020; He et al., 2023) to speed up the training. However, they solely focus on response collection time (Liu et al., 2022a; b) and overlook the time required to acquire adequate resources. Additionally, optimizing individual FL job performance is insufficient as the deployment scale of FL applications continues to grow.
FL Resource Manager. Large companies including Apple, Meta and Google have proposed their FL infrastructures; however, FL resource management is not their primary focus. These resource managers simply adopt random device-to-job matching in various forms, resulting in suboptimal scheduling delays and response collection times.
@@ 7 Conclusion
In this paper, we introduce our FL resource manager, Venn, to enable efficient sharing of a large amount of heterogeneous device resources among multiple FL jobs with diverse requirements. Venn incorporates a contention-aware job scheduling algorithm and a resource-aware device-to-job matching algorithm to minimize the average JCT for FL jobs. Our evaluation over various real-world FL workloads shows that Venn achieves up to 1.88X improvement on average JCT.
Appendix A ILP Formulation of IRS
@@
Appendix B Theoretical Insight to the Heuristic of IRS
@@
Given a diverse set of FL jobs with one round request, if jobs are scheduled optimally in terms of the average JCT, first within each job group and then across job groups, the resulting average JCT is optimal.
@@
Appendix C Effectiveness of Venn Scheduling Heuristic
@@
As shown in Figure 14(a), there are $m_{A}$ requests that ask for devices with 1GB memory and $m_{B}$ jobs that request devices with 2GB memory, resulting in two job groups $A$ and $B$. The devices constantly check-in and execute one FL task, where 100% devices with memory size $\geq$ 1GB and xof the devices have memory size $\geq$ 2GB. Note that, the proof is not limited to the contention pattern draw in Figure 14(a), it can be generalized to job group with intersected resource contention and give the same conclusion.
@@
where $m_{A}^{\prime}$, $m_{B}^{\prime}$ represents the number of remaining jobs whose queuing delay may be affected by this prioritization. Since the future resource allocation is set by the previous initial allocation or assumed to be given, $m_{A}^{\prime}$, $m_{B}^{\prime}$ are feasible to get. We prioritize the first job from Group A only if $\Delta t<0$, which gives $\frac{m_{A}^{\prime}}{1-x}>\frac{m_{B}^{\prime}}{x}$, otherwise we stick with the original plan. $\Delta t<0$ is actually the prototype of the scheduling decision as in Algorithm 1 line 13.
@@
P33fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ 5 Evaluation
  • Venn reduces average job completion time (JCT) by up to 1.88×1.$8$ compared to state-of-the-art baselines, without compromising model accuracy, across diverse real-world CL workloads (§5.2).
  • Venn outperforms its counterparts by intelligently scheduling jobs and matching devices to jobs, leveraging different design components (§5.3).
  • Venn’s benefits are robust under a wide range of CL workloads and environment setups (§5.5).
@@ 5.1 Experiment Setup
To rigorously evaluate Venn, we employ a two-pronged approach. First, we have developed a high-fidelity simulator that replays client and job traces, effectively emulating the dynamics of the scheduling environment for large-scale evaluations. Second, we deploy real CL systems to execute actual CL jobs at a smaller scale of devices.
Resources.
@@
To faithfully emulate heterogeneous device runtimes, networking, and availability, we use device traces from FedScale Lai et al. (2021a) (Figure 2) and AI Benchmark Ignatov et al. (2019) (Figure 8(a)). Each unique device trace is limited to one CL job per day for realism.
CL jobs.
@@
We evaluate Venn over multiple synchronous CL jobs Yang et al. (2018), where each successful training round requires a minimum of 80% target participants to report back within a deadline, which is set to be 5min - 15min depending on the round demand. Our approach also extends to asynchronous CL jobs, as scheduling decisions depend solely on remaining resource demand, not request submission timing. Jobs are drawn from diverse real-world CL applications Xu et al. (2023); Yang et al. (2018); Hard et al. (2019); Ramaswamy et al. (2019); Warden (2018); Sadilek et al. (2021), whose resource demand is depicted in Figure 8(b). In the real CL experiment, each job aims to train a ResNet-18 He et al. (2016) and MobileNet-V2 Sandler et al. (2018) on FEMNIST dataset.
Workloads.
@@
Our evaluation includes five workload scenarios that sample differently from the job trace in Figure 8(b) to rigorously evaluate Venn’s performance. Even: Sampled from all jobs, which is the default trace. Small: Uniformly sampled only from jobs with below-average total demand. Large: Uniformly sampled only from jobs with above-average total demand. Low: Uniformly sampled only from jobs with below-average demand per round. High: Uniformly sampled only from jobs with above-average demand per round. Default simulation and real-world workloads contain 50 and 20 jobs, respectively. Jobs arrive via a Poisson process with a 30-min average inter-arrival.
@@
Device requirements are stratified into four categories based on the CPU and memory capacities (Figure 8(a)) to create various resource contention patterns where the eligible resources for each job may overlap, contain, or be within the eligible resources of other jobs. By default, each job is randomly mapped to one category among General resources, Compute-Rich resources, Memory-Rich resources, High-Performance resources.
Baselines.
@@
We compare Venn against FIFO, SRSF, and an optimized random matching. The random matching algorithm typically assigns devices to eligible jobs randomly but is optimized here to schedule jobs in a randomized order, reducing round abortions under contention and establishing a stronger baseline. Note that we only run Venn with the starvation prevention strategy in Section 5.5.
Metrics.
@@
Our primary performance metrics include the average job completion time (JCT). Note that while Venn does not explicitly optimize for CL job accuracy, it does not adversely affect it either.
@@
Table 1: Summary of improvements on average JCT over random matching on different CL workloads.
@@ 5.2 End-to-End Performance
Venn achieves better average JCT Improvement. We assess the performance of different scheduling algorithms by evaluating its performance over different workloads. We report the average JCT speed-up for each scheduling algorithm compared to the random scheduling in Table 1. We observe that our scheduling algorithm consistently provides stable improvements in the average JCT across various workloads, which underscores the robustness of Venn.
Venn achieves faster convergence without affecting accuracy. We report the final model test accuracy of CL jobs under different schedules with the help of our CL system at a smaller experiment scale. As shown in Figure 10, we observe that Venn does not affect the final model test accuracy but speeds up the overall convergence process.
Venn has negligible overhead. We emulated a large number of CL jobs and groups to evaluate the scheduler’s scalability. Our results in Figure 10 demonstrate that the latency incurred by one-time triggering for scheduling and matching remains low, even with a substantial increase in job and group numbers. This is due to its low time complexity of $\max(O(m\log m),O(n^{2}))$, where $m$ is the number of jobs and $n$ is the number of device groups.
@@ 5.3 Performance Breakdown
Table 3: Breakdown of average JCT improvement across jobs that ask for General resources, Compute-rich resources, Memory-rich resources and High-performance resources. Venn benefits more on jobs that ask for scarcer resources.
@@
To analyze the impact of Venn on different types of jobs, we break down jobs by their total demands and device requirements (Figure 8(a)), and then analyze the average JCT improvement for each type. Table 2 and Table 3 quantifies how Venn improves average JCT across varying total demands (25th, 50th, 75th percentiles) and eligibility types (General, Compute-rich, Memory rich, High-performance ). Notably, jobs with smaller total demands and scarcer resources benefit the most from Venn.
5.4 Case Study on Biased Workload
@@ 5.5 Ablation Study
Impact of number of jobs.
We evaluate Venn with different numbers of CL jobs arriving over time. As the number of jobs increases, resource contention becomes more pronounced, highlighting the importance of efficient scheduling under such conditions. We present the average JCT speed-up with different numbers of jobs in even workload to demonstrate the effectiveness of our algorithm. As shown in Figure 13, Venn consistently provides improvement across various numbers of jobs, with its benefits becoming more pronounced as the number of jobs increases.
Impact of number of tiers.
@@
Impact of fairness knob.
We incorporated a fairness knob ($\epsilon$) to strike a balance between performance and fairness. We tune the value of $\epsilon$ and report the average JCT speed-up against these values in Figure 14(a). The results demonstrate that, as $\epsilon$ increases, the JCT speed-up tends to decrease. As shown in Figure 14(b), the percentage of jobs that meet the fair-share JCT increases with the $\epsilon$, where ϵ=italic-$\epsilon=$2 gives 69% of jobs receive their fair-share JCT. This observation highlights the trade-off between performance and fairness within our CL resource scheduling algorithm, which can be fine-tuned by adjusting the value of $\epsilon$.
6 Related Work
Cluster resource managers. There are many cluster resource managers that schedule resources with constraints Thinakaran et al. (2017); Ghodsi et al. (2013); Narayanan et al. (2020). Existing ML cluster resource managers mainly focus on managing the stable data-center resources like GPU and CPU (Xiao et al., 2018; Mohan et al., 2022) in order to improve JCT, utilization and fairness (Peng et al., 2018; Chaudhary et al., 2020). Some research delves into GPU-specific optimizations Gu et al. (2019); Yu et al. (2021); Hwang et al. (2021), while others co-design resource managers with the specific characteristics of ML workloads Qiao et al. (2021); You et al. (2023). However, they are mostly designed for data center and fail to capture the level of availability and heterogeneity of CL resources, nor do they consider both scheduling delay and response collection time of CL jobs.
CL client selectors. Several recent works have studied client selection at the single CL job level. In addition to enforce device requirements including software version, hardware capacity and data quality, they further cherry-pick clients based on their state, system and statistical utility (Lai et al., 2021b; Abdelmoniem et al., 2023; Balakrishnan et al., 2022; Jee Cho et al., 2022; Chai et al., 2020; He et al., 2023) to speed up the training. However, they solely focus on response collection time (Liu et al., 2022a; b) and overlook the time required to acquire adequate resources. Additionally, optimizing individual CL job performance is insufficient as the deployment scale of CL applications continues to grow.
CL resource managers. Large companies including Apple, Meta and Google have proposed their CL infrastructures; however, CL resource management is not their primary focus. These resource managers simply adopt random device-to-job matching in various forms, resulting in suboptimal scheduling delays and response collection times.
@@ 7 Conclusion
In this paper, we present Venn, our CL resource manager designed to efficiently share large-scale heterogeneous device resources among multiple CL jobs with diverse requirements. Venn features a contention-aware job scheduling algorithm and a resource-aware device-to-job matching algorithm, aiming to minimize the average job completion time (JCT) for CL workloads. Our evaluation using a range of real-world CL workloads demonstrates that Venn improves average JCT by up to 1.88×1.$8$.
Acknowledgements
We thank the anonymous reviewers for valuable feedback. This work was supported in part by NSF grant CNS-2106184 and a grant from Cisco.
Appendix A Detailed Venn Responsibility
Venn delegates responsibilities such as device selection, device fault tolerance, and privacy protection to individual CL jobs. Device failures are both inevitable and difficult to predict in CL. Rather than imposing a one-size-fits-all solution, Venn empowers CL jobs to take the reins on fault tolerance based on their specific workloads and objectives. Therefore, Venn offloads handling device fault tolerance to CL jobs, who can better detect and react to device failures (e.g., deciding the amount of overcommit Bonawitz et al. (2019)). Similarly, Venn offers CL jobs the freedom to design their own device selectors Lai et al. (2021b), where they can incorporate customized resource criteria into their requests. Venn also does not interfere with job-specific privacy solutions such as secure aggregation Bonawitz et al. (2016); Huba et al. (2022) or differential privacy Geyer et al. (2017); Xu et al. (2023).
Appendix B ILP Formulation of IRS
@@
Appendix C Theoretical Insight to the Heuristic of IRS
@@
Given a diverse set of CL jobs with one round request, if jobs are scheduled optimally in terms of the average JCT, first within each job group and then across job groups, the resulting average JCT is optimal.
@@
Appendix D Effectiveness of Venn Scheduling Heuristic
@@
As shown in Figure 15(a), there are $m_{A}$ requests that ask for devices with 1GB memory and $m_{B}$ jobs that request devices with 2GB memory, resulting in two job groups $A$ and $B$. The devices constantly check-in and execute one CL task, where 100% devices with memory size $\geq$ 1GB and xof the devices have memory size $\geq$ 2GB. Note that, the proof is not limited to the contention pattern draw in Figure 15(a), it can be generalized to job group with intersected resource contention and give the same conclusion.
@@
where $m_{A}^{\prime}$, $m_{B}^{\prime}$ represents the number of remaining jobs whose queuing delay may be affected by this prioritization. Since the future resource allocation is set by the previous initial allocation or assumed to be given, $m_{A}^{\prime}$, $m_{B}^{\prime}$ are feasible to get. We prioritize the first job from Group A only if $\Delta t<0$, which gives $\frac{m_{A}^{\prime}}{1-x}>\frac{m_{B}^{\prime}}{x}$, otherwise we stick with the original plan. $\Delta t<0$ is actually the prototype of the scheduling decision as in Algorithm 1 line 15.